2x - y =6
x + 2y =8

Answer by jim_thompson5910(35256)

Hint:

2x - y =6
x + 2y =8


2(2x - y) =2*6
x + 2y =8


4x - 2y = 12
x + 2y = 8
------------
5x+0y = 20

5x = 20

Solve that for x. Then use that solution for x to get the solution for y.